# A Volumetric Approach to Privacy of Dynamical Systems

### v3

@@ 4.2 Randomization
${\alpha}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\right],$
and
${\beta}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\right].$
Since $\left(x_{k}-\underline{X}_{k|k-1}\right)$ is the radius between the actual public state $X_{k}=x_{k}$ and the lower endpoint of $\mathcal{X}_{k|k-1}$, and $\left(\overline{X}_{k|k-1}-x_{k}\right)$ is the radius between $x_{k}$ and the upper endpoint of $\mathcal{X}_{k|k-1}$, the random set ${\mathcal{S}}_{k|k}^{x}$ is a subset of $\mathcal{X}_{k|k-1}$ that contains the actual public state. Also, we have $\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ as
$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)$
$=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
@@
We next restrict $\mathcal{S}^{x}_{k|k}$ be the subset of the observation set $\mathcal{M}^{x}_{k|k}$, and optimize $\mathcal{M}^{x}_{k|k}$ to enhance the privacy level.
@@ 4.3 Privacy Filter Optimization
[…]
$\mathbf{P_{1}}:$ $\max_{\mathcal{M}^{x}_{k|k}}\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ (42)
[…]
where $\mathcal{S}^{x}_{k|k}\subseteq\mathcal{M}^{x}_{k|k}$ is required by the structure of the privacy filter, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ describes the prior space of $\mathcal{M}^{x}_{k|k}$, and $\mathrm{Vol}\left({\mathcal{M}}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ enforces the requirement of data utility. We next demonstrate that the optimization problem $\mathbf{P_{1}}$ can be solved via linear programming.
@@
$\mathbf{P_{2}}:\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$ (48)
@@
$\left\{\begin{array}[]{c}p_{k-1|k}^{\Delta z}\geqslant 0\\ p_{k-1|k}^{\Delta z}\geqslant p_{k-1|k-1}^{z}-p_{k-1|k}^{m,z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\\ \end{array}\right.,\text{for}\;Z=X,Y$ (59)
where $\epsilon^{y}\geqslant 0$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{R}^{2n_{x}}$ and $p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}\in\mathcal{R}^{n_{x}}$.
Proof 4.1.
@@
Figure 3: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.01$.
Figure 4: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.5$.
As a result, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3. Although the linear programming problem $\mathbf{P_{2}}$ is deterministic, we can demonstrate that the adversary cannot reduce its uncertainty by reversing the optimization process.
Proposition 8.
If the adversary infers the private state via the inference attack (2)-(32), then it cannot find a smaller interval subset of $\mathcal{M}_{k|k}^{x,\star}$ containing the actual public state $X_{k}=x_{k}$ via reverting the linear programming.
Proof 4.2.
On the one hand, the actual public state $X_{k}=x_{k}$ can be any element of $\mathcal{X}_{k|k-1}$, and $S_{k}^{x}$ is a random subset of $\mathcal{X}_{k|k-1}$ containing $x_{k}$. Thus, the upper or lower endpoint of $S_{k}^{x}$ may coincide with $x_{k}$.
On the other hand, since $\mathcal{M}_{k|k}^{x,\star}$ is the optimal observation set obtained from the linear programming, for any $S_{k}^{x}$ in the vicinity of $\mathcal{M}_{k|k}^{x,\star}$, i.e., $S_{k}^{x}\in\left\{S_{k}^{x}\left|\lim_{\left\|\Delta\right\|_{1}\rightarrow 0}{S_{k}^{x}+\Delta=\mathcal{M}_{k|k}^{x,\star}}\right.\right\}$, the optimal solution remains $\mathcal{M}_{k|k}^{x,\star}$.
As a result, it is possible that the upper or lower endpoint of $\mathcal{M}_{k|k}^{x,\star}$ corresponds to the actual public state. Thus, the minimal interval observation set for the adversary is $\mathcal{M}_{k|k}^{x,\star}$.
According to Proposition 8, even if the adversary knows the structure of the privacy filter, it cannot revert the linear programming to obtain a smaller interval containing $x_{k}$ to reduce uncertainty. In other words, the privacy filer is robust against the reverse optimization attack.
@@ 5 Numerical Verification
$\left(\mathcal{W}_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.74&1.91&1.94&2.01\\ \end{matrix}\right],$
$\left(\mathcal{W}_{k}^{y}\right)^{\top}=\left[\begin{matrix}0.91&0.23&0.95&0.43\\ \end{matrix}\right].$
@@
In our simulation, the initial states are uniformly sampled from the bounded sets (60)-(61). To simulate the approximate periodic fluctuations in demand and productivity, the actual process noises are set to be
@@
We first plot the trajectories of system states and their interval tubes in Fig.3 and Fig.4 under the optimal privacy filter design for different values of $\epsilon^{x}$. We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$. The pink areas in these figures represent the uncertainty sets of system states. As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate. However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s estimation of the production rate to become less accurate. Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
@@
We analyze the utility-privacy trade-off performance of the optimal privacy filtering policy and compare it with the truncated Gaussian mechanism, which is proposed in [18] for differential privacy. In the truncated Gaussian mechanism, illustrated in Fig. 5, the original state $x_{k}$ is perturbed by additive noise $v_{k}$ drawn from a truncated Gaussian distribution defined over the interval $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, with zero mean and variance $\left(\epsilon^{x}\right)^{2}$. Subsequently, the perturbed observation set is released as follows:
$\mathcal{M}_{k|k}^{x}=\left[\begin{array}[]{c}z_{k}-\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ z_{k}+\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ \end{array}\right].$ (62)
Since the noise $v_{k}$ is constrained within $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, the released public state $x_{k}$ is guaranteed to reside within the interval $\mathcal{M}_{k|k}^{x}$, and the utility of $x_{k}$ is quantified by $\epsilon^{x}$.
The privacy-utility trade-off is evaluated by plotting the privacy level of the production rate against the utility of the inventory under different approaches, as depicted in Fig. 6. The results indicate that increasing the utility of inventory leads to a reduction in the privacy level of the production rate, highlighting the inherent trade-off between data utility and privacy protection. Compared to the truncated Gaussian mechanism, the optimal privacy filter achieves a higher privacy level while maintaining lower data distortion.
Figure 6: The privacy level of the private state and utility of the public state.
@@ 6 Conclusion
In this paper, we developed a volumetric privacy measure for dynamic systems with UBB noise. We defined the inference attack that an adversary uses to estimate the private state and introduced a volumetric measure to evaluate the privacy level. We provided computational approaches based on interval analysis and analyzed the theoretical properties of the proposed measure. Furthermore, we proposed an optimization-based approach for privacy filter design to defend the system against inference attacks. The effectiveness of our method was demonstrated through a production-inventory case study.
@@ Appendix A Proof of Lemma 2
At the time step $k=0$, the adversary only has prior knowledge, i.e., $\mathcal{Y}_{0|-1}$, therefore, its inference set is $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$.
@@
$\mathcal{S}=\Psi\left(A\right)\mathcal{M}+\Psi\left(B\right)\mathcal{W},$ (63)
where
$\Psi\left(N\right)=\left[\begin{matrix}\frac{N+\left|N\right|}{2}&\frac{N-\left|N\right|}{2}\\ \frac{N-\left|N\right|}{2}&\frac{N+\left|N\right|}{2}\\ \end{matrix}\right].$ (64)
Also, we can compute the radius and center of the intervl $\mathcal{S}$ via
$p^{s}=\left|A\right|p^{m}+\left|B\right|p^{w},$ (65)
$c^{s}=Ac^{m}+Bc^{w}.$ (66)
When $A_{1}$ and $A_{2}$ are invertible, we have
$X_{k-1}=A_{1}^{-1}X_{k}+\left(-A_{1}^{-1}A_{2}\right)Y_{k-1}+\left(-A_{1}^{-1}B_{1}\right)W_{k}^{x},$
and
$Y_{k-1}=A_{2}^{-1}X_{k}+\left(-A_{2}^{-1}A_{1}\right)X_{k-1}+\left(-A_{2}^{-1}B_{1}\right)W_{k}^{x}.$
According to Lemma 9, the tightest intervals for (2.3) and (2.3) are (2) and (2). Then, the intersection of different intervals, i.e., (12) and (13), can be computed with (24) and (27). Finally, the one-step forward reachable set (14) can be approximated with the tightest interval (32) based on Lemma 9, and the calibrated uncertainty set $\mathcal{X}_{k|k}$ and the tightest prior inference set $\mathcal{Y}_{k|k-1}$ can be approximated similarly. Finally, since at the time step $k=0$, the backward calibration (2.3) and (2.3) is not available, the adversary can only calibrate the public state set with its prior knowledge $\mathcal{X}_{0|-1}$ and the observation set $\mathcal{M}_{0|0}^{x}$ according to (36).
@@ Appendix B Proof of Lemma 3
$p_{k-1|k}^{m,y}=\left|A_{2}^{-1}\right|p_{k|k}^{m,x}+\left|A_{2}^{-1}A_{1}\right|p_{k-1|k-1}^{x}+\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x},$
@@
$p_{k|k}^{y}=\left|A_{3}\right|p_{k-1|k}^{x}+\left|A_{4}\right|p_{k-1|k}^{y}+\left|B_{2}\right|p_{k}^{w,y}.$
Therefore, with (B), we have
@@ Appendix C Proof of Theorem 5
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}=\Phi\left(A_{3}\right)\Delta\mathcal{X}_{k-1|k}+\Phi\left(A_{4}\right)\Delta\mathcal{Y}_{k-1|k},$
@@
$\Delta\mathcal{X}_{k-1|k}$ $=\mathcal{X}_{k-1|k-1}-\mathcal{M}_{k-1|k}^{x}$
@@
$\Delta\mathcal{Y}_{k-1|k}$ $=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$
@@
$p_{k|k}^{\Delta y}=\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y},$ (73)
@@
$=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}+\max\left\{\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},0\right\}$
@@
$\left.\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\right\},\text{for}\,Z=X,Y,$ (74)
@@
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
$=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
=$\overset{(a)}{=}\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
$\leqslant\left\|A_{3}\right\|\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
$=\left\|A_{3}\right\|\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$ (75)
@@
Besides, given an interval $\mathcal{X}$, we can express it with its center point and radius, i.e., $\mathcal{X}=\left[\begin{array}[]{c}\frac{c-p}{2}\\ \frac{c+p}{2}\end{array}\right]$. Therefore, we have (C),
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
$=$ $\left\|\overline{Y}_{k|k-1}-\overline{Y}_{k|k}\right\|_{1}+\left\|\underline{Y}_{k|k-1}-\underline{Y}_{k|k}\right\|_{1}$
$=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
@@
$\geqslant$ $2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$ (76)

### v4

@@ 4.2 Randomization
${\alpha}_{k}\!\in\!\left[\!0,\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\right]\!\!,{\beta}_{k}\!\in\!\left[\!0,\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\right]\!\!.$
Since $\left(x_{k}-\underline{X}_{k|k-1}\right)$ is the radius from the actual public state $X_{k}=x_{k}$ to the lower endpoint of $\mathcal{X}_{k|k-1}$, and $\left(\overline{X}_{k|k-1}-x_{k}\right)$ is the radius from $x_{k}$ to the upper endpoint of $\mathcal{X}_{k|k-1}$, the random set ${\mathcal{S}}_{k|k}^{x}$ becomes a subset of $\mathcal{X}_{k|k-1}$ that contains the actual public state. Also, we can shown $\mathcal{S}^{x}_{k|k}$ satisfies the utility constraint as follows,
$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
@@
We next restrict $\mathcal{S}^{x}_{k|k}$ be the subset of the observation set $\mathcal{M}^{x}_{k|k}$, and optimize $\mathcal{M}^{x}_{k|k}$ to improve the privacy level.
@@ 4.3 Privacy Filter Optimization
In this subsetion, we demonstrate that the optimization problem $\mathbf{P_{1}}$ based on the interval inference can be solved via linear programming.
@@
$\mathbf{P_{2}}:$ $\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$
@@
$\left\{\begin{array}[]{c}p_{k-1|k}^{\Delta z}\geqslant 0\\ p_{k-1|k}^{\Delta z}\geqslant p_{k-1|k-1}^{z}-p_{k-1|k}^{m,z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\\ \end{array}\right$., (63)
where $p_{k-1|k}^{\Delta x}\!\in\mathcal{R}^{n_{x}}$, $Z=X,Y$, $\epsilon^{y}\geqslant 0$ and $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{R}^{2n_{x}}$.
Proof 4.2.
@@
Figure 3: Inference attack results after applying the optimal privacy filter: (a), (b) Estimated $X_{k}$ and (e), (f) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.01$; (c), (d) Estimated $X_{k}$ and (g), (h) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.5$.
Consequently, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3.
@@ 5 Numerical Verification
$\left(\!\mathcal{W}_{k}^{x}\!\right)^{\top}\!\!=\!\left[\!\begin{matrix}1.74&1.91&1.94&2.01\\ \end{matrix}\!\right]\!\!,\left(\!\mathcal{W}_{k}^{y}\!\right)^{\top}\!\!=\!\left[\!\begin{matrix}0.91&0.23&0.95&0.43\\ \end{matrix}\!\right]\!\!.$
@@
In our simulation, the initial states are uniformly sampled from the bounded sets (64)-(65). To simulate the approximate periodic fluctuations in demand and productivity, the actual disturbance are set to be
@@
We first plot the trajectories of system states and their interval tubes in Fig.3 under the optimal privacy filter design for different values of $\epsilon^{x}$. We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$. The pink areas in these figures represent the uncertainty sets of system states. As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate. However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s central estimation of the production rate to become less accurate. This observation numerically verifies Theorem 9, confirming that a higher privacy level prevents the adversary from refining its incorrect central estimation. Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
@@
We also evaluate the utility–privacy trade-off achieved by the proposed optimal privacy-filtering policy and compare it with two benchmark mechanisms: the noiseless quantization method presented in [21] and the truncated Gaussian mechanism for differential privacy introduced in [31]. In the quantization-based approach, the state $x_{k}$ is processed through a static quantizer that satisfies the utility constraint, and the quantization bin containing $x_{k}$ is publicly released as $\mathcal{M}_{k|k}^{x}$. In contrast, the truncated Gaussian mechanism, illustrated in Fig. 4, perturbs the original state $x_{k}$ with additive noise $v_{k}$ drawn from a zero-mean truncated Gaussian distribution supported on the interval $\left[-\epsilon^{x}/2,\,\epsilon^{x}/2\right]$ and having variance $\left(\epsilon^{x}\right)^{2}$. The perturbed observation set is then released in the form
$\mathcal{M}_{k|k}^{x}=\left[\begin{array}[]{c}z_{k}-\frac{\epsilon^{x}}{2n_{x}}\cdot\mathbf{1}_{n_{x}}\\ z_{k}+\frac{\epsilon^{x}}{2n_{x}}\cdot\mathbf{1}_{n_{x}}\end{array}\right],$
where $n_{x}$ denotes the dimension of $x_{k}$ and $\mathbf{1}_{n_{x}}$ is the all-ones vector of length $n_{x}$. Because the additive noise $v_{k}$ is bounded within $\left[-\epsilon^{x}/2,\,\epsilon^{x}/2\right]$, the publicly released state is guaranteed to lie within the interval $\mathcal{M}_{k|k}^{x}$.
The privacy-utility trade-off was evaluated by plotting the average privacy level of the production rate against the average utility of the inventory over 100 random trajectories with horizon 100, as illustrated in Fig. 5. For a fair and clear comparison, the privacy level and utility values for the truncated Gaussian mechanism are normalized to the range $[0,1]$, and its same scaling parameters are applied to the other two mechanisms. The results demonstrate that an increase in inventory utility corresponds to a reduction in the privacy level of the production rate, thereby confirming the intrinsic trade-off between data utility and privacy protection. Besides, the quantization mechanism and the truncated Gaussian mechanism have similar performance in privacy-utility trade-off. Note that these two mechanisms assume static domains of states, while the sets of states in systems subject to disturbance can be estimated and described more precise as addressed by the inference attack. The proposed volumetric approach estimates the time-varying private state set and adjusts the output accordingly.
Furthermore, as discussed in [31], if the volume (e.g., interval length) of domain satisfies certain conditions, the truncated Gaussian mechanism ensures differential privacy. However, given the volume constraint, the shape of the public state set can be arbitrary, and certain shapes may lead to substantial volumetric leakage of the private state after inference attacks. The proposed volumetric method explicitly accounts for this by considering the geometry of the set based on the assumed inference attack, not only its volume. Therefore, it achieves higher privacy levels while maintaining lower data distortion compared with the two static mechanisms.
It is worth noting that an adversary could employ more sophisticated estimation techniques, e.g., CCG-based approximation, to infer the private set more accurately from the interval observations provided by the privacy filter. Nevertheless, as shown in Fig. 6, the proposed privacy filter still outperforms the other two mechanisms, leveraging knowledge of the underlying state evolution to reduce conservativeness.
Figure 5: Interval-based inference given the interval privacy filter: the privacy level of the private state and utility of the public state.
Figure 6: CCG-based inference given the interval privacy filter: the privacy level of the private state and utility of the public state.
@@ 6 Conclusion
In this paper, we develop a volumetric framework for privacy analysis and defense in dynamic systems subject to bounded disturbance. An inference attack, whereby an adversary estimates the private state, is formalized, and a volumetric measure is introduced to quantify the resulting privacy level. We develop computational methods based on interval analysis, and establish the theoretical properties of the measure. Furthermore, we propose an optimization-based approach for privacy filter design to defend the system against inference attacks. The effectiveness of our method is demonstrated through a production-inventory case study.
It is noted that the performance of the volumetric privacy measure inherently depends on the selected set-membership estimation techniques, and its evaluation accuracy varies with different set representations. Future research will focus on developing approximation methods that ensure improved accuracy, robustness and broader applicability.
@@ Appendix A Proof of Lemma 6
At the time step $k=0$, the adversary only has prior knowledge, i.e., $\mathcal{Y}_{0|-1}$, therefore, its inference set is $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$. Also, since at the time step $k=0$, the backward calibration (6) and (7) is not available, the adversary can only calibrate the public state set with its prior knowledge $\mathcal{X}_{0|-1}$ and the observation set $\mathcal{M}_{0|0}^{x}$ according to (40).
@@
$\mathcal{S}=\Psi\left(A\right)\mathcal{M}\oplus\Psi\left(B\right)\mathcal{W}.$
Also, we can compute its radius and center via
$p^{s}=\left|A\right|p^{m}$ $+\left|B\right|p^{w},$
$c^{s}=Ac^{m}$ $+Bc^{w}.$
According to Lemma 13, the tightest intervals for (6) and (7) are (24) and (25). Then, the intersection of different intervals, i.e., (8) and (9), can be computed with (28) and (31). Finally, the one-step forward reachable set (10) can be approximated with the tightest interval (36) based on Lemma 13, and the calibrated uncertainty set $\mathcal{X}_{k|k}$ and the tightest prior inference set $\mathcal{Y}_{k|k-1}$ can be approximated similarly.
@@ Appendix B Proof of Lemma 8
$p_{k-1|k}^{m,y}\!=\!\left|\!A_{2}^{-1}\!\right|p_{k|k}^{m,x}\!+\!\left|\!A_{2}^{\!-\!1}\!A_{1}\!\right|p_{k\!-\!1|k-\!1}^{x}\!+\!\left|\!A_{2}^{\!-\!1}\!B_{1}\!\right|p_{k}^{w,x}\!,\!\!$ (66)
@@
$p_{k|k}^{y}=\left|A_{3}\right|p_{k-1|k}^{x}+\left|A_{4}\right|p_{k-1|k}^{y}+\left|B_{2}\right|p_{k}^{w,y}.$ (67)
By substituting (66) and $p_{k-1|k}^{y}\leqslant p_{k-1|k}^{m,y}$ into (67), we have
@@ Appendix C Proof of Theorem 9
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k}=\Phi\left(A_{3}\right)\Delta\mathcal{X}_{k-1|k}\oplus\Phi\left(A_{4}\right)\Delta\mathcal{Y}_{k-1|k},$
@@
$\Delta\mathcal{X}_{k-1|k}$ $=\mathcal{X}_{k-1|k-1}\setminus\mathcal{X}_{k-1|k}$
@@
$\Delta\mathcal{Y}_{k-1|k}$ $=\mathcal{Y}_{k-1|k-1}\setminus\mathcal{Y}_{k-1|k}$
@@
$p_{k|k}^{\Delta y}=\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y},$
@@
$\!\!\!\!\overline{Z}_{k-\!1|k-\!1}\!-\!\overline{M}_{k-\!1|k}^{z},\underline{M}_{k-1|k}^{z}\!-\!\underline{Z}_{k-\!1|k-\!1}\left.\right\},\text{for}\,Z=X,Y,\!\!$ (74)
@@
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ $=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
$\overset{(a)}{\leqslant}\left\|A_{3}\right\|_{1}\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|_{1}\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
$=\left\|A_{3}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$
@@
Besides, given an interval $\mathcal{X}$, we can express it with its center point and radius, i.e., $\mathcal{X}=\left[\begin{array}[]{c}\frac{c-p}{2}\\ \frac{c+p}{2}\end{array}\right]$. Therefore, we have
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
@@
$\geqslant$ $2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$